OpenStudy (anonymous):

a salesman earned $900 last week. He has a retainer of $270 and earns a commision of 12% of sales over $2000. Find the value of sales.

OpenStudy (anonymous):

$270 + 12% ( S - $2,000 ) = $900 -> .12S = 900 - 270 + 240 => S = 87,000 / 12 = $7,250